### v2

# references.bib

Abstract
We propose the Block Coordinate Descent Network Simplex (BCDNS) method for solving large-scale discrete Optimal Transport (OT) problems. BCDNS integrates the Network Simplex (NS) algorithm with a block coordinate descent (BCD) strategy, decomposing the full problem into smaller subproblems per iteration and reusing basis variables to ensure feasibility. We prove that BCDNS terminates in a finite number of iterations with an exact optimal solution, and we characterize its per-iteration complexity as $O(sN)$, where $s\in(0,1)$ is a user-defined parameter and $N$ is the total number of variables. Numerical experiments demonstrate that BCDNS matches the classical NS method in solution accuracy, reduces memory footprint compared to Sinkhorn algorithm, achieves speed-ups of up to tens of times over the classical NS method, and exhibits runtime comparable to high-precision Sinkhorn algorithm.
Keywords: Optimal Transport, Network Simplex, Block Coordinate Descent, Sinkhorn, Wasserstein Distance
1.   Introduction
The Optimal Transport (OT) problem is to find the most efficient way to transport resources between a finite number of locations while minimizing transportation costs. It is rooted in mathematical optimization and has a rich history dating back to the 18th century. Gaspard Monge [monge1781deblais] first proposed it in 1781. Then in the mid-20th century, Leonid Kantorovich [kantorovich1942translocation] extended Monge’s framework to encompass continuous distributions, reformulating it as a linear programming model and significantly broadening its theoretical foundation and practical applications. In this thesis, we focus specifically on solving the discrete OT problem.
The OT problem has found applications in various disciplines. In machine learning, the OT problem has been used for tasks such as domain adaptation [courty2016optimal] and generative modeling  [arjovsky2017wasserstein]. In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [galichon2016optimal]. In biology, the OT problem has been applied to analyze genetic expression patterns [schiebinger2019optimal].
The OT problem has found applications in various disciplines. In machine learning, the OT problem has been used for tasks such as domain adaptation [courty2016optimal] and generative modeling [arjovsky2017wasserstein]. In economics, the OT problem has provided insights into resource allocation, market equilibrium, and pricing mechanisms [galichon2016optimal]. In biology, the OT problem has been applied to analyze genetic expression patterns [schiebinger2019optimal]. In geophysics and signal processing, the OT problem has been employed for seismic waveform comparison and inversion. [liao2022fast].
Several methods have been developed to solve the OT problem. Each method has distinct strengths and limitations. The Hungarian method is a classical approach for solving linear assignment problems and can also be applied to OT problems [kuhn1955hungarian]. It achieves exact optimal solutions by constructing a cost matrix and iteratively reducing reducing dimensions of the matrix. However, its time complexity is $O(N^{2})$, where $N$ is the number of variables in OT problem, and hence it is not applicable for the large-scale OT problem. The Sinkhorn algorithm is a popular method for the large-scale OT problem due to its computational efficiency [cuturi2013sinkhorn]. It solves the OT problems with regularization of entropy, which is an approximation of the OT problem. Therefore it provides approximate solutions of the OT problem, and it may not be suitable for applications that require exact solutions. In recent years, many improvements based on the Sinkhorn algorithm have been proposed, such as the Greenkhorn algorithm [altschuler2017Near] and the Adaptive Primal–Dual Accelerated Gradient Descent (APDAGD) algorithm [dvurechensky2018Computational]. However, these methods still operate within the Sinkhorn framework and do not overcome the limitation of producing only approximate OT solutions. In contrast, the Network Simplex (NS) method, a specialized variant of the simplex method, is well known to yield exact solutions. Michael D. Grigoriadis [grigoriadis1986efficient] reported that the NS method is reliable among several algorithms. However, the NS method is not a polynomial time algorithm and it takes $O(N)$ time in each iteration, which leads to slow computational performance when applied to large-scale problems. We note that the above three methods needs at least $O(N)$ space complexity. To overcome the difficulty of large space complexity for the OT problem, Yue Xie et al. [xie2023randomized] proposed the Block Coordinate Descent (BCD) method for solving the OT problem. Their method constructs a small subproblem by randomly selecting variables based on predefined rules and solves it in each iteration. Since the subproblem is small, the BCD method does not require the O(N) space complexity. However, since the BCD method proposed by Yue Xie et al. adapts the random variable selection, it is not guaranteed to find an optimal solution within a finite number of iterations. Moreover, it sometimes takes a large number of iterations to get a reasonable approximate solution.
The motivation behind this research lies in the growing demand for efficient methods capable of solving large-scale OT problems with highly accurate solutions. As data sizes continue to grow, we recognize the increasing importance of solving large-scale OT problems with both speed and accuracy. To address these challenges, we aim to develop a method that enables fast computation of exact solutions for large-scale OT problems.
To achieve the goal of efficiently solving large-scale OT problems, we introduce the idea of the BCD method. The BCD method divides a large-scale OT problem into smaller subproblems, enabling a reduction in computational complexity. Although the BCD method is generally considered slow when exact solutions are required, we observe that the optimal solutions of the OT problem are typically sparse—due to the relatively small number of equality constraints compared to the number of variables—which significantly improves the efficiency of BCD. By employing the NS method to solve the subproblems, we guarantee that the final solution is also exact.
In this paper, we propose the Block Coordinate Descent Network Simplex (BCDNS) method, a novel algorithm that integrates the BCD and NS methods to achieve computational efficiency and solution accuracy. Our approach reduces computational complexity by decomposing large-scale OT problems into smaller subproblems, while the NS method ensures that each subproblem is solved with precision. Numerical experiments demonstrate significant performance gains of the BCDNS method over traditional approaches, addressing key limitations of existing methods.
In the BCD method, the selection of the block coordinate is crucial. Compared to existing BCD method [xie2023randomized] that randomly selects variables based on matrix properties, we propose that the BCDNS method first selects the basis variables from the previous subproblem as the initial solution for the next subproblem. This strategy ensures that the subproblem always has a feasible and an optimal solution. Moreover, by integrating the principles of the NS method, we ensure that the BCDNS method converges to the exact optimal solution within a finite number of computations. Furthermore, we categorize variables into two groups based on their likelihood of becoming optimal basis variables. This targeted selection process prioritizes variables with a higher probability of being optimal basis variables, thereby improving the efficiency of finding the optimal solution.
The outline of this paper is as follows. In Section 2, we introduce the foundational concepts of the OT problem and review several existing methods for solving it. In Section 3, we describe the core principles of the proposed BCDNS method. We first present its prototype along with the corresponding theoretical proofs. Then we develop a concrete implementation based on the prototype. Section 4 reports numerical experiments, including parameter estimation for the BCDNS method and a comparative performance analysis against existing approaches. Finally, Section 5 concludes the paper by summarizing the key contributions and discussing potential directions for future research.
2.   Discrete Optimal Transport
We consider a discrete OT problem [villani2009optimal] between $n$ supply locations and $m$ demand locations. Let $\{1,2,\dots,n\}$ denote the index set of supply points, and $\{1,2,\dots,m\}$ denote that of demand points.
Each supply point $i\in\{1,\dots,n\}$ provides a non-negative amount of mass $p_{i}\geq 0$, and each demand point $j\in\{1,\dots,m\}$ requires an amount $q_{j}\geq 0$. The total supply and demand are assumed to be balanced:
$\sum_{i=1}^{n}p_{i}=\sum_{j=1}^{m}q_{j}.$
Let $C=\{c_{ij}\}\in\mathbb{R}^{n\times m}$ be the cost matrix, where each entry $c_{ij}\geq 0$ denotes the cost of transporting one unit of mass from supply point $i$ to demand point $j$. A transport plan is represented by a matrix $X=\{x_{ij}\}\in\mathbb{R}^{n\times m}$, where $x_{ij}$ is the amount of mass transported from $i$ to $j$.
The discrete OT problem can then be formulated as the following linear program:
$\min_{X\in\mathbb{R}_{\geq 0}^{n\times m}}\quad$ $\sum_{i=1}^{n}\sum_{j=1}^{m}c_{ij}x_{ij}$ (1)
subject to $\sum_{j=1}^{m}x_{ij}=p_{i},\quad\forall i=1,\dots,n,$
$\sum_{i=1}^{n}x_{ij}=q_{j},\quad\forall j=1,\dots,m,$
$x_{ij}\geq 0,\quad\forall i=1,\dots,n,\;j=1,\dots,m.$
To express this in standard matrix-vector form, we define the following notation. Let $N=nm$ be the total number of transport variables, and let $M=n+m$ be the number of equality constraints. We vectorize the matrix $X$ into a column vector $\mathbf{x}\in\mathbb{R}^{N}$ by stacking its columns, i.e.,
$\mathbf{x}=\operatorname{vec}(X)=[x_{11},x_{21},\dots,x_{n1},x_{12},\dots,x_{nm}]^{\top}.$
We also define:
  • the cost vector $\mathbf{c}=\operatorname{vec}(C)\in\mathbb{R}^{N}$,
  • the supply vector $\mathbf{p}=[p_{1},\dots,p_{n}]^{\top}\in\mathbb{R}^{n}$,
  • the demand vector $\mathbf{q}=[q_{1},\dots,q_{m}]^{\top}\in\mathbb{R}^{m}$,
  • the constraint vector $\mathbf{b}=[\mathbf{p}^{\top},\,-\mathbf{q}^{\top}]^{\top}\in\mathbb{R}^{M}$.
  • The constraint matrix
    $A=\begin{bmatrix}I_{n}&I_{n}&\cdots&I_{n}\\ -1_{n}^{\top}&0&\cdots&0\\ 0&-1_{n}^{\top}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-1_{n}^{\top}\end{bmatrix}\in\mathbb{R}^{M\times N},$
    where $I_{n}$ denotes the $n\times n$ identity matrix, and $1_{n}$ is an $n$-dimensional column vector of ones.
With this notation, we can rewrite the OT problem as:
$\min_{\mathbf{x}\in\mathbb{R}^{N}}\quad$ $\mathbf{c}^{\top}\mathbf{x}$ (2)
subject to $A\mathbf{x}=\mathbf{b},$
$\mathbf{x}\geq 0.$
3.   Existing Methods for Optimal Transport
In this section, we review several representative algorithms for solving the discrete OT problem. These include the classical NS method, the Sinkhorn algorithm based on entropy regularization which is widely used for large-scale problems, and the recently proposed BCD method.
3.1 Network Simplex Method for Optimal Transport
As shown in Appendix A, the constraint matrix $A\in\mathbb{R}^{M\times N}$ in the OT problem has rank $M-1$. This allows us to partition the set of variable indices $\mathcal{N}=\{1,\dots,N\}$ into two subsets: a basis index set $\mathcal{B}\subseteq\mathcal{N}$ with $|\mathcal{B}|=M-1$, and its complement $\mathcal{N}\setminus\mathcal{B}$. The columns of $A$ corresponding to $\mathcal{B}$, denoted $A_{B}\in\mathbb{R}^{M\times(M-1)}$, form a full-rank submatrix. The basic variables $\mathbf{x}_{B}$ can then be uniquely determined by the system $A_{B}\mathbf{x}_{B}=\mathbf{b}$, while the non-basic variables $\mathbf{x}_{N}$ are fixed at zero. A basic feasible solution further satisfies the non-negativity constraint $\mathbf{x}_{B}\geq 0$.
To determine whether a given basic solution is optimal, we examine its reduced-cost vector. Define the dual potential vector $\boldsymbol{\pi}\in\mathbb{R}^{M}$ via the dual feasibility condition $\boldsymbol{\pi}=(A_{B}^{\top})^{-1}\mathbf{c}_{B}$, where $\mathbf{c}_{B}$ is the cost subvector corresponding to the basis $B$. Then, for any index $k\in N\setminus B$, the reduced cost is computed as $r_{k}=c_{k}-\boldsymbol{\pi}^{\top}A_{k}$. Due to the structure of the OT problem, each column $A_{ij}$ has exactly one $+1$ and one $-1$, so the reduced cost simplifies to the closed-form expression
$r_{ij}=c_{ij}-(\pi_{i}-\pi_{n+j}).$
Optimality of the current basic solution is certified if all reduced costs are non-negative, i.e., $r_{ij}\geq 0$ for all $i,j$. If a variable $x_{i^{*}j^{*}}$ has $r_{i^{*}j^{*}}<0$, it indicates a direction of descent in the objective, and that variable is introduced into the basis.
To perform a pivot, we add the index $(i^{*},j^{*})$ to the basis $B$, inducing a unique cycle in the associated flow graph. Then determine the step size $\Delta$ by the minimum flow along the negative edges of the cycle, ensuring feasibility. Subsequently, update the flow: $x_{i^{*}j^{*}}\leftarrow\Delta$, and for each edge $(u,v)$ along the cycle, the value of $x_{uv}$ is incremented or decremented by $\Delta$ depending on the cycle direction. Finally, the variable that reaches zero and remove it from the basis, completing the pivot step [vanderbei2001linear].
Each iteration of the NS method involves computing the dual potential vector $\boldsymbol{\pi}$ in $O(M)$ time, evaluating all reduced costs in $O(N)$, and performing pivoting operations in $O(M)$. Since each pivot strictly improves the objective function by at least $-r_{i^{*}j^{*}}>0$, and the number of possible basic feasible solutions is finite, the NS method terminates in a finite number of iterations.
3.2 Block Coordinate Descent Method for Optimal Transport
In this subsection we introduce the BCD method for the OT problem in the LP form (2).
The BCD method iteratively selects a subset of coordinates $\mathcal{H}\subseteq\mathcal{N}=\{1,2,\dots,N\}$. For any feasible solution $\mathbf{x}$, we define a subproblem of the BCD method
$\𝒩D(\mathbf{x};\mathcal{H})\triangleq\arg\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}.$ (3)
$\𝒩q(\mathbf{x};\mathcal{H})\triangleq\min_{\mathbf{d}\in\mathbb{R}^{N}}\left\{\mathbf{c}^{T}(\mathbf{x}+\mathbf{d})\mid\mathbf{x}+\mathbf{d}\geq 0,\;A\mathbf{d}=0,\;d_{i}=0,\;\forall i\in\mathcal{N}\backslash\mathcal{H}\right\}.$ (4)
Here, $D(\mathbf{x};\mathcal{H})$ represents the set of optimal solutions for the subproblem with $\mathcal{H}$, and $q(\mathbf{x};\mathcal{H})$ denotes the optimal objective value corresponding to $d\in D(\mathbf{x};\mathcal{H})$ [tseng2009block].
The iterative BCD algorithm proceeds as follows:
  • Initialization: Start with a feasible solution $\mathbf{x_{0}}$ and an initial block $\mathcal{H}_{0}\subseteq\{1,\dots,N\}$.
  • Subproblem Optimization: For the current solution $\mathbf{x^{k}}$, solve the subproblem restricted to the working set $\mathcal{H}_{k}$ to find the descent direction $\mathbf{d^{k}}\in D(\mathbf{x^{k}};\mathcal{H}_{k})$. Update the solution as:
    $\mathbf{x^{k+1}}=\mathbf{x^{k}}+\mathbf{d^{k}}.$
  • Optimality Check: If optimality is achieved, terminate the algorithm. Else adjust the block $\mathcal{H}_{k}$ for the next iteration using predefined rules or adaptive heuristics.
The existing BCD method [xie2023randomized], randomly selects block coordinates based on certain properties of specific matrices. In [xie2023randomized], the authors propose the Accelerated Random Block Coordinate Descent algorithm and establish the following convergence theorem.
Theorem 1.
The gap between the current objective function value and the optimal objective function value converges with rate $1-v$, i.e.,
$\mathbf{c}^{\top}\mathbf{x}^{k+1}-\mathbf{c}^{\top}\mathbf{x}^{*}\leq(1-v)(\mathbf{c}^{\top}\mathbf{x}^{k}-\mathbf{c}^{\top}\mathbf{x}^{*}).$
Here, $v$ satisfies
$v\geq\zeta\frac{n(\tilde{p}-2)}{(n^{2}-3)(n!)^{2}},$
where $\zeta$ is a predefined probability parameter in the range $(0,1)$, and $\tilde{p}$ is a parameter representing the properties of the matrix in the range $[3,n]$.
These approaches provide a simple mechanism for constructing subproblems by selecting block coordinates. However, when the size n of the OT problems is large, $v$ becomes extremely small, which leads to slow convergence. Furthermore, these approaches do not account for the sparsity of the solution of the OT problem. Many variables that ultimately take zero values continue to be included in the block coordinates at each iteration. Moreover, these approaches have an inherent limitation: they cannot guarantee convergence to the optimal solution within a finite number of iterations. This limitation arises from the possibility of extreme cases where the randomly selected block of coordinates rare includes the optimal basis variables.
3.3 Sinkhorn Algorithm for Optimal Transport
Instead of solving the classical OT problem defined in (1), the Sinkhorn algorithm [cuturi2013sinkhorn] introduces an entropy regularization term
$\min_{X\in\mathbb{R}_{+}^{n\times m}}\;\sum_{i=1}^{n}\sum_{j=1}^{m}c_{ij}x_{ij}-\varepsilon H(X),$
where $\varepsilon>0$ is the regularization parameter, and $H(X)$ is the entropy function defined as
$H(X)=-\sum_{i=1}^{n}\sum_{j=1}^{m}x_{ij}\log x_{ij}.$
The parameter $\varepsilon$ controls the strength of the entropy regularization, balancing the trade-off between the classical transport cost and the entropy of the transport plan.
Introduce Lagrange multipliers $\alpha=(\alpha_{i})_{i=1}^{n}$ and $\beta=(\beta_{j})_{j=1}^{m}$ for the supply and demand constraints, respectively. The Lagrangian function of the regularized OT problem is
$\mathcal{L}(X,\alpha,\beta)=\sum_{i=1}^{n}\sum_{j=1}^{m}c_{ij}x_{ij}-\varepsilon H(X)+\sum_{i=1}^{n}\alpha_{i}\left(p_{i}-\sum_{j=1}^{m}x_{ij}\right)+\sum_{j=1}^{m}\beta_{j}\left(q_{j}-\sum_{i=1}^{n}x_{ij}\right).$
Taking the partial derivative of $\mathcal{L}$ with respect to $x_{ij}$ and setting it to zero, we obtain the optimality condition
$\frac{\partial\mathcal{L}}{\partial x_{ij}}=c_{ij}+\varepsilon(\log x_{ij}+1)-\alpha_{i}-\beta_{j}=0.$
Solving for $x_{ij}$, we find
$x_{ij}=\exp\left(\frac{\alpha_{i}+\beta_{j}-c_{ij}}{\varepsilon}-1\right).$
Let $u_{i}=e^{\alpha_{i}/\varepsilon}$ and $v_{j}=e^{\beta_{j}/\varepsilon}$, then the optimal transport plan can be expressed as
$x_{ij}=u_{i}e^{-c_{ij}/\varepsilon}v_{j}.$
Define $K=e^{-C/\varepsilon}$ as the matrix obtained by element-wise exponentiating the cost matrix $C=\{c_{ij}\}$. The transport plan $X$ can then be written as:
$X=\text{diag}(\mathbf{u})K\text{diag}(\mathbf{v}),$
where $\text{diag}(\mathbf{u})$ and $\text{diag}(\mathbf{v})$ are diagonal matrices whose diagonal entries are the elements of the vectors $\mathbf{u}$ and $\mathbf{v}$, respectively.
From the equality constraints $X\mathbf{1}_{m}=\mathbf{p}$ and $X^{\top}\mathbf{1}_{n}=\mathbf{q}$, we derive the following update rules
$\mathbf{u}^{(t+1)}=\frac{\mathbf{p}}{K\mathbf{v}^{(t)}},$
$\mathbf{v}^{(t+1)}=\frac{\mathbf{q}}{K^{\top}\mathbf{u}^{(t+1)}}.$
Here, the division is performed element-wise.
The Sinkhorn algorithm terminates when the change in the scaling factors $\mathbf{u}$ and $\mathbf{v}$ becomes sufficiently small
$\|\mathbf{u}^{(t)}-\mathbf{u}^{(t-1)}\|<\delta,\quad\|\mathbf{v}^{(t)}-\mathbf{v}^{(t-1)}\|<\delta,$
where $\delta>0$ is a small predefined tolerance.
The regularization parameter $\varepsilon$ introduces a smoothing effect in the optimal transport problem. As $\varepsilon$ decreases, the solution becomes closer to the exact OT plan, but the algorithm may converge more slowly. Conversely, larger values of $\varepsilon$ result in faster computations, but with a higher approximation error in the transport plan.
4.   Block Coordinate Network Simplex Methods with Basis Variable Succession
The goal of the NS method is to identify the optimal basis variables through iterative pivoting. While each NS iteration efficiently updates the basis structure, computing reduced costs for all variables incurs a cost of $O(N)$ per iteration.
To reduce this per-iteration cost, we adopt the BCD strategy: at each iteration, we restrict attention to a selected subset $\mathcal{H}_{k}\subseteq\mathcal{N}$ of variables and solve a corresponding subproblem using the NS method.
The OT problem is particularly well-suited to such block decomposition. In general LPs with $M$ constraints, computing the reduced cost for a single variable requires $O(M)$ operations. However, due to the special network structure of the OT problem, each reduced cost can be computed in $O(1)$ time via the relation
$r_{ij}=c_{ij}-(\pi_{i}-\pi_{n+j}),$
We now summarize the notation and assumptions used throughout this section:
  • $\mathbf{x}^{k}\in\mathbb{R}^{N}$: the basic feasible solution at iteration $k$,
  • $\mathbf{x}^{*}$:the optimal solution to the full OT problem (2),
  • $\mathcal{B}_{k}\subseteq\{1,\dots,N\}$: the set of basis indices at iteration $k$,
  • $\mathcal{B}^{\ast}\subseteq\{1,\dots,N\}$: the basis of a globally optimal solution,
  • $\mathcal{H}_{k}\subseteq\{1,\dots,N\}$: the block of variable indices selected at iteration $k$,
  • $\bar{c}^{k}:=\bar{c}(\mathbf{x}^{k},\mathcal{B}_{k})$: the reduced cost vector at $\mathbf{x}^{k}$.
4.1 Basis Variables Succession
In BCD schemes for the OT problem, the convergence behaviour hinges on the construction of the working set $\mathcal{H}_{k}$ at each iteration $k$.
The following theorem identifies the precise conditions under which a restricted NS step on a block $\mathcal{H}_{k}$ yields a globally optimal solution. It also shows when global optimality cannot be achieved.
Theorem 2.
Apply the network–simplex method to the subproblem restricted to $\mathcal{H}_{k}$, and set
$d^{k}\in\mathcal{D}\bigl{(}\mathbf{x}^{k};\mathcal{H}_{k}\bigr{)},\qquad\mathbf{x}^{k+1}=\mathbf{x}^{k}+d^{k}.$
Then:
  1. (i)
    If $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$ and $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the iterate $\mathbf{x}^{k+1}$ is globally optimal.
  2. (ii)
    If $\mathcal{B}_{k}\not\subseteq\mathcal{H}_{k}$, then—even when $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$—the iterate $\mathbf{x}^{k+1}$ cannot be globally optimal.
Proof.
(i) Because $\mathcal{B}_{k}\subseteq\mathcal{H}_{k}$, the NS subproblem starts with basis $\mathcal{B}_{k}$. Pivoting within $\mathcal{H}_{k}$ yields a basis ^$\widehat{\mathcal{B}}\subseteq\mathcal{H}_{k}$ satisfying
[…]
Since $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$, the subproblem contains an optimal basis of the full problem; NS therefore stops only when no negative reduced costs remain in $\mathcal{H}_{k}$. As dual variables depend solely on the basis,
^$\widehat{\mathcal{B}}\cap\mathcal{H}_{k}=\mathcal{B}^{\ast},$
so condition (5) extends to all indices $i=1,\dots,N$:
[…]
By complementary slackness, $\mathbf{x}^{k+1}$ is globally optimal.
(ii) Assume $j\in\mathcal{B}_{k}\setminus\mathcal{H}_{k}$ with $\mathbf{x}^{k}_{j}>0$. Because $j\notin\mathcal{H}_{k}$, the subproblem cannot pivot $j$ or alter its value, hence
$\mathbf{x}^{k+1}_{j}=\mathbf{x}^{k}_{j}>0.$
In any optimal OT solution all non–basic variables are zero. Since $j\notin\mathcal{B}^{\ast}$ (because $\mathcal{B}^{\ast}\subseteq\mathcal{H}_{k}$), optimality would require $x^{\ast}_{j}=0$, contradicting $\mathbf{x}^{k+1}_{j}>0$. Therefore $\mathbf{x}^{k+1}$ cannot be optimal. ∎
The above analysis motivates a practical strategy, which we call basis variable succession: at each iteration, we selecte the block $\mathcal{H}_{k}$ includes the basis variables $\mathcal{B}_{k}$ obtained from the previous subproblem. This design contrasts with existing BCD approaches [xie2023randomized] that select blocks randomly.
we can solve the subproblem in iteration $k$ with complexity $O(L_{k}|\mathcal{H}_{k}|)$, which is significantly lower than the $O(LN)$ complexity of solving the full problem directly, where $L_{k}$ and $L$ denote the number of NS iterations for the subproblem and the full problem, respectively.
This strategy offers two practical advantages:
  1. 1.
    The subproblem always includes all constraints and at least $M-1$ variables, ensuring feasibility.
  2. 2.
    Reusing the previous basis variables as the initialization eliminates the need for recomputing a starting basis in the NS method.
4.2 A Prototype of Block Coordinate Descent Network Simplex Methods and its Finite Termination
Following the strategy of §4.1, choose a block
$\mathcal{H}_{k}\supseteq\mathcal{B}_{k},\qquad\lvert\mathcal{H}_{k}\rvert\geq M-1,$
and split it into
$\mathcal{B}_{k},\qquad\mathcal{I}_{k}=\mathcal{H}_{k}\setminus\mathcal{B}_{k}.$
If $\mathcal{I}_{k}$ is picked arbitrarily, the subproblem
$\mathcal{D}\!\bigl{(}\mathbf{x}^{k};\mathcal{H}_{k}\bigr{)}$
might admit only the zero direction, yielding the same basis $\mathcal{B}_{k+1}=\mathcal{B}_{k}$ and $\mathbf{x}^{k}=\mathbf{x}^{k+1}$.
Define the index set of negative reduced costs
$\mathcal{N}_{k}\;=\;\bigl{\{}\,\ell\in\{1,\dots,nm\}:\;\bar{c}^{\,k}_{\ell}<0\bigr{\}}.$
From iteration $k$ on we impose the rule
$\mathcal{I}_{k}\cap\mathcal{N}_{k}\;\neq\;\varnothing.$ (6)
If $\mathcal{N}_{k}=\varnothing$, then $\bar{c}^{\,k}\geq 0$ and, by complementary slackness, $\mathbf{x}^{k}$ is already globally optimal.
Under condition (6) each subproblem produces a strictly improving search direction, ensuring monotone decrease of the objective value. The next theorem establishes finite termination and global optimality of the resulting sequence $\{\mathbf{x}^{k}\}$.
Theorem 3.
The proposed prototype of BCDNS method satisfies the following properties:
  1. 1.
    Monotonicity and Convergence: The objective function decreases monotonically at each iteration, and the method converges to the global optimal solution.
  2. 2.
    Finite Termination: The method terminates within a finite number of iterations.
Proof of Part 1.
Monotonicity: For any iteration $k$, based on the selection rule for $\mathcal{I}_{k}$, there exists at least one $l\in\mathcal{I}_{k}$ such that the reduced cost of $x_{l}$ is negative. By the properties of the reduced cost, adding $x_{l}$ to the basis variables decreases the total cost, i.e.,
$q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\{l\})<q(\mathbf{x}^{k};\mathcal{B}_{k}).$
where $q(\mathbf{x}^{k};\mathcal{I}_{k})$ is defined by (4) as the optimal value of the subproblem associated with the index set $\mathcal{I}_{k}$. Since $l\in\mathcal{I}_{k}$, it follows that:
$q(\mathbf{x}^{k};\mathcal{H}_{k})=q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\mathcal{I}_{k})\leq q(\mathbf{x}^{k};\mathcal{B}_{k}\cup\{l\})<q(\mathbf{x}^{k};\mathcal{B}_{k}).$
Furthermore, since $\mathcal{J}_{k}$ is the optimal basis variables of the previous subproblem, we have:
$q(\mathbf{x}^{k};\mathcal{B}_{k})=q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$
Combining the above, it follows that:
$q(\mathbf{x}^{k};\mathcal{H}_{k})<q(\mathbf{x}^{k-1};\mathcal{H}_{k-1}).$
Thus, the cost function decreases monotonically at each iteration.
Global Optimality: After solving each subproblem $\mathcal{D}(\mathbf{x}^{k};\mathcal{H}_{k})$, we compute the reduced costs of all variables. If all reduced costs satisfy $r_{ij}\geq 0$, the algorithm terminates; otherwise, the next iteration begins. By the properties of the reduced cost, this condition guarantees that the final solution is globally optimal. ∎
Proof of Part 2.
The total number of possible combinations of the basis variables among $N$ variables is at most $\binom{N}{M-1}$. Based on the selection rule, in each iteration, the set $\mathcal{I}_{k}$ always contains at least one variable with a negative reduced cost. Denote the variable with the smallest reduced cost as $x_{l}$. When solving the subproblem using the NS method, $x_{l}$ will be added to the basis variables, while one of the current basis variables $\mathcal{B}_{k}$will leave the basis. This implies that each iteration examines at least one new combination of the basis variables. Consequently, after at most $\binom{N}{M-1}$ iterations, all possible combinations of the basis variables will have been examined. Therefore, the algorithm is guaranteed to terminate within a finite number of iterations. ∎
4.3 Grouped Selection Block Coordinate Descent Network Simplex method
Since an optimal solution to the OT problem is typically sparse, most variables take value zero and do not belong to the optimal basis. Consequently, selecting $\mathcal{I}_{k}$ uniformly at random often results in the inclusion of irrelevant variables, slowing convergence.
To address this issue, we propose a grouped block coordinate selection strategy that partitions the variable index set $\{1,\dots,N\}$ into two disjoint subsets:
  • $\mathcal{R}$: indices of variables deemed likely to enter the optimal basis $\mathcal{B}^{*}$,
  • $\mathcal{S}$: indices of variables with low likelihood of being part of the optimal basis $\mathcal{B}^{*}$.
Let $0<s<t\leq 1$ be fixed sampling parameters. At each iteration, the reduced costs are computed for a randomly selected subset of size $tN$ from $\mathcal{R}$, together with a small subset from $\mathcal{S}$ for exploration. Among these, we select the $sN$ variables with the smallest reduced costs to form the block $\mathcal{I}_{k}$, as these variables are most promising for descent directions.
If all selected variables have non-negative reduced costs, the algorithm infers that none are useful for further improvement and moves them to set $\mathcal{S}$. The selection process is then repeated until condition (6) is satisfied.
This strategy guides the search toward structurally relevant directions and improves computational efficiency. Based on this mechanism, we introduce the Grouped Selection Block Coordinate Descent Network Simplex (GS-BCDNS) method. The full algorithm is presented below.
Algorithm 1 Grouped Selection Block Coordinate Descent Network Simplex Method (GS-BCDNS)
Initialization:
Find an initial feasible solution $X^{0}\in\mathbb{R}^{n\times m}$. Set the initial basis variables as $\mathcal{J}_{0}$. Choose parameters $s,t\in(0,1)$ such that $s<t$. Initialize $\mathcal{S}=\emptyset$ and =$\mathcal{R}=\{1,2,\ldots,N\}$. Shuffle $\mathcal{R}$ randomly. Define the threshold sequence $\{e_{k}\}$.
for $k=0,1,2,\ldots$ do
     Step 1: Select the first $\min(|\mathcal{R}|,tN)$ variables from $\mathcal{R}$ and $\min(|\mathcal{S}|,0.1tN)$ variables from $\mathcal{S}$. Compute the reduced costs for all selected variables and sort them in ascending order. Select the $sN$ variables with smallest reduced costs to form $\mathcal{I}_{k}$. If all selected variables have non-negative reduced costs, move them to $\mathcal{S}$ and repeat Step 1.
     Step 2: Move the non-selected variables with positive reduced costs to $\mathcal{S}$, and return variables with non-positive reduced costs to $\mathcal{R}$.
     Step 3: Set $\mathcal{H}_{k}=\mathcal{I}_{k}\cup\mathcal{B}_{k}$.
     Step 4: Use the NS method to solve the subproblem $\mathcal{D}(x^{k};\mathcal{H}_{k})$ and obtain $d^{k}\in\mathcal{D}(x^{k};\mathcal{H}_{k})$. Update the basis to $\mathcal{B}_{k+1}$. Move variables with reduced cost $r_{ij}<e_{k}$ to $\mathcal{R}$, and move the remaining variables to $\mathcal{S}$.
     Step 5: Update $x^{k+1}:=x^{k}+d^{k}$.
     Step 6: Optimality check: If $\mathcal{R}=\emptyset$, compute reduced costs for all variables. If $r_{ij}\geq 0$ for all $i,j$, terminate; otherwise, return to Step 1.
end for
Let $\tilde{k}$ denote the total number of outer iterations, and let $l_{k}$ represent the number of NS iterations required to solve the subproblem at iteration $k$. The time complexity of the GS-BCDNS method is given by:
$O\left(\tilde{k}tN+\sum_{k=1}^{\tilde{k}}l_{k}sN\right).$
The complexity clearly depends on the choice of sampling parameters $s$ and $t$. Decreasing their values reduces the cost per iteration, but increases the total number of iterations $\tilde{k}$. Thus, careful tuning of $s$ and $t$ is essential to balance convergence speed and efficiency.
In each iteration, the space complexity of the GS-BCDNS method is:
$O(|\mathcal{H}_{k}|)=O(sN+M-1).$
Compared to the Sinkhorn algorithm, which requires $O(N)$ space, the GS-BCDNS method is more memory-efficient. However, relative to existing BCD methods, the memory usage depends on the size of the selected block $\mathcal{H}_{k}$ [xie2023randomized].
5.   Numerical Experiments
In this section, we perform numerical experiments on various instances of OT problems1.11footnotetext: 1 Numerical experiments were conducted on a computer with an Intel Core i7-12700H CPU (2.70 GHz) and 16 GB of RAM. The implementation was performed using Python 3.8, and the code has been made publicly available on GitHub at https://github.com/842422492/BCDNS.git. We first solve randomly generated OT problems using the GS-BCDNS method to evaluate the impact of parameters $s$ and $t$ on algorithmic efficiency. Subsequently, we compute the Wasserstein distance between uniform and normal distributions using different algorithms to compare their performance.
5.1 Experiment Settings
For simplicity, we assume that the numbers of supply and demand nodes are equal, i.e., $n=m$.
Test Problem 1: Ten randomly generated discrete OT problems using randomly sampled numbers. $n=1000$.
Test Problem 2: The Wasserstein distance between a one-dimensional uniform distribution on $[-1,1]$ and a standard normal distribution on $\mathbb{R}$. $n=1000$.
Test Problem 3: The Wasserstein distance between a two-dimensional uniform distribution on $[-1,1]^{2}$ and a standard normal distribution on $\mathbb{R}^{2}$. $n=1000$.
Test Problem 4: The Wasserstein distance between a three-dimensional uniform distribution on $[-1,1]^{3}$ and a standard normal distribution on $\mathbb{R}^{3}$. $n=1000$.
Test Problem 5: The Wasserstein distance between a one-dimensional uniform distribution on $[-1,1]$ and a standard normal distribution for different problem sizes. $n$ starts from 50 and increases by 50 up to 300.
The discrete OT problems in Test Problem 1 were generated as follows: two sets of $n=1000$ random numbers were sampled from a uniform distribution on $[1,1000]$ to represent the supply and demand distributions $P$ and $Q$, respectively. Then adjuste the two sets to ensure that their sums were equal. Additionally, a cost matrix of size $1000\times 1000$ was generated from the same uniform distribution.
The Wasserstein distances in Test Problem 2, 3, and 4 were estimated as follows: $n$ samples were randomly drawn from the uniform and normal distributions, denoted as $y$ and $\tilde{y}$, respectively. For each pair, i.i.d. samples $\{\tilde{y}_{i}\}$ and $\{y_{j}\}$, $1\leq i,j\leq 1000$, were used to approximate the two continuous probability measures. The cost matrix was defined as: $c_{i,j}=\text{dist}(\tilde{y}_{i},y_{j})^{2},$ and the probability mass vectors were set as $\mathbf{p}=\mathbf{q}=(1/n,\dots,1/n)^{T}\in\mathbb{R}^{n}$.
Figure 1: The point distributions in Test Problem 2, 3 and 4
Algorithm Settings:
  • Sinkhorn (Stabilized) Algorithm: As used in [xie2023randomized], we follow the method proposed in [cuturi2013sinkhorn] for computing the Wasserstein distance. Let $\gamma$ be the entropy regularization coefficient, which is set to $\gamma=1/(4\log n)$ as suggested in [dvurechensky2018computational]. We consider the parameter settings $\varepsilon=10^{-4},10^{-3},0.01,0.1$. Sinkhorn iterations are projected onto the feasible region using a rounding procedure, specifically Algorithm 2 in [altschuler2017near]. To mitigate numerical instability, all updates are performed in log space using the LogSumExp function. The algorithm terminates when either the change in the objective function value falls below $\delta=10^{-6}$ or the runtime exceeds 2000 seconds.
  • GS-BCDNS: When solving Test Problem 1, the parameter $s$ was varied from $1/n$ to 0.9, while the parameter $t$ was varied from $2/n$ to 1. For Test Problems 2, 3, 4 and 5, the values of $s$ and $t$ were fixed at $5/n$ and $10/n$, respectively. For each Test Problem, set $\{e_{k}\}$ to positive infinity and calculate the initial basic feasible solution using the Northwest Corner Rule.
5.2 Effect of Parameters $s$ and $t$
In the BCDNS method, the parameters $s$ and $t$ play a crucial role. Here, we vary the parameters $s$ and $t$ to solve Test Problem 1. The number of reduced cost computations serves as a key indicator in the BCDNS method. For each case, we record the corresponding number of pivots and reduced cost computations. And then we calculate the averages of these metrics to evaluate the algorithm’s performance. In this context, the number of pivots refers to the number of iterations performed by the NS method.
Figure 2: Effects of $s$ and $t$ on the number of pivots and reduced cost computations
Figure 1 illustrates the effects of $s$ and $t$ on two key performance metrics: the number of pivots and the total number of reduced cost computations. We can observe that the number of pivots decreases as $s$ increases, as larger subproblem sizes reduce the total iterations required by the NS method. However, as $s$ increases, the size of each subproblem also grows, which, while reducing the number of pivots, increases the computational cost of each pivot. Similarly, increasing $t$ enhances the quality of the variables selected for the subproblem, further reducing the number of pivots.
For small values of $s$, increasing $s$ reduces the total number of reduced cost computations by enabling more efficient solutions to larger subproblems. However, beyond a certain threshold, further increases in $s$ lead to an overall rise in computational effort due to the greater complexity of individual subproblems. The effect of $t$ is less pronounced, except when $s$ is very small, in which case increasing $t$ significantly increases the computational burden, as the algorithm approaches the behavior of the standard NS method. Extensive numerical experiments consistently reveal that reduced cost computations are minimized when $t$ is approximately $2s$.
5.3 Comparison with the Existing Methods
In this chapter, we primarily evaluate the performance differences among the GS-BCDNS method, the Sinkhorn algorithm, and the NS method.
Figure 3: Comparison of algorithms to computer Wasserstein distance
As described in Section 4.1, we apply the Sinkhorn algorithm with different values of $\epsilon$ and the GS-BCDNS method to solve Test Problems 2, 3 and 4. The stopping criterion for the Sinkhorn algorithm is set such that the algorithm terminates when the change in the objective function value falls below $\delta=10^{-6}$ or when the runtime exceeds 2000 seconds. The GS-BCDNS method is tested with fixed parameters $s=5/n$ and $t=10/n$. In the evaluation, the x-axis represents the runtime, while the y-axis represents the gap between the current objective function value and the optimal objective function value, where the optimal solution is obtained using the interior-point method implemented in Python’s SciPy library.
The results in Figure 3 indicate that for larger values of $\epsilon$ (i.e, $\epsilon=0.1$ and $\epsilon=0.01$), the Sinkhorn algorithm exhibits a faster convergence speed than the GS-BCDNS method; however, the approximation error remains relatively large. When $\epsilon=0.001$, the convergence speed of the Sinkhorn algorithm is comparable to that of the GS-BCDNS method, but its final convergence is slightly slower. For extremely small values of $\epsilon$, the convergence speed of the Sinkhorn algorithm is significantly slower than that of the GS-BCDNS method.
Figure 4: Comparison of algorithms to solve Test Problem 5
Figure 4 presents the results of solving Test Problem 5 using both the NS method and the GS-BCDNS method. The y-axis of Figure 4 represents the ratio of the NS method’s runtime to that of the GS-BCDNS method. The results demonstrate that the GS-BCDNS method significantly outperforms the NS method across all tested datasets.
6.   Conclusion
In this paper, we review the existing methods for solving the OT problems, including the NS method, the BCD method, and the Sinkhorn algorithm. Then we propose the BCDNS methods and explain thier theoretical foundations and the basis variable succession strategy. Furthermore, we provide a detailed algorithmic framework and established theoretical guarantees for finite termination and convergence to the exact optimal solution.
Extensive numerical experiments demonstrate that the BCDNS method achieves superior performance in terms of runtime and computational cost compared to the NS method. While the BCDNS method converges slightly slower than the Sinkhorn algorithm with large values of $\epsilon$, it has the advantage of guaranteeing convergence to an exact solution. Moreover, compared to the Sinkhorn algorithm with small values of $\epsilon$, the BCDNS method exhibits even faster convergence. This makes the BCDNS method particularly valuable for large-scale OT problems that require exact solutions.
Future work could explore several directions. First, improving the selection strategy for block coordinates may further enhance convergence speed. Second, extending the BCDNS method to more general network flow problems could broaden its applicability.
Appendix A Proof of $\operatorname{rank}(A)=M-1$
Proof.
Let
$A=\begin{bmatrix}I_{n}&I_{n}&\cdots&I_{n}\\ -\mathbf{1}_{n}^{\top}&0&\cdots&0\\ 0&-\mathbf{1}_{n}^{\top}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-\mathbf{1}_{n}^{\top}\end{bmatrix}\in\mathbb{R}^{M\times N},\quad M=n+m,\;N=nm,$
where $I_{n}$ is the $n\times n$ identity matrix and $\mathbf{1}_{n}$ is the $n$-vector of all ones.
(i) $\operatorname{rank}(A)\leq M-1$.
Denote the rows of $A$ by $r_{1},\dots,r_{M}$. Observe that
[…]
because each of the first $n$ rows contributes exactly one “$+1$” in each block and each of the last $m$ rows contributes exactly one “$-1$”, so they cancel out. Hence the rows are linearly dependent, and
$\operatorname{rank}(A)\;\leq\;M-1.$
(ii) $\operatorname{rank}(A)\geq M-1$.
Interpret $A$ as the node–arc incidence matrix of the complete bipartite directed graph $K_{n,m}$ with $M=n+m$ nodes and $nm$ arcs. Choose any spanning tree $T$ of this graph, which has exactly $M-1$ edges. Let
$B\;=\;A(:,E(T))\;\in\;\mathbb{R}^{M\times(M-1)}$
be the submatrix of $A$ whose columns correspond to the edges in $T$. Since $T$ is a tree, its incidence matrix has rank $M-1$, and in particular there is an $(M-1)\times(M-1)$ submatrix of $B$ with determinant ±$\pm 1$. Therefore
$\operatorname{rank}(A)\;\geq\;M-1.$
Combining (i) and (ii) gives
$\operatorname{rank}(A)\;=\;M-1.$
[…]